OpenStudy (anonymous):

y=f(t)=-11+120t-16t^2 ft.. A) when, if ever, is the projectile 189 ft above the ground?

OpenStudy (saifoo.khan):

189 =-11+120t-16t^2 Solve.

OpenStudy (anonymous):

oki

OpenStudy (anonymous):

200=8t(15-2t)

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

wht do i next?

OpenStudy (saifoo.khan):

25 = t(15 - 2t) 25 = 15t - 2t^2 Quadratic.

OpenStudy (anonymous):

so solve for t?

OpenStudy (anonymous):

so i got t=1.4, -8.9